# Exploring the operational limits of Mirnov coil arrays in stellarators by means of a synthetic diagnostic

### v1

@@ VI Conclusions and future work
Finally, solving the wave equation by Fourier methods and including the response of an homogeneous plasma through its cold dielectric tensor, we have analyzed the plasma response for modes with $n=0$ and arbitrary mode number $m$. Although the model is limited in several important aspects, all of them discussed in the text, it shows that measured mode numbers may differ from the real ones up to $\Delta m=2$. Combining this result with the one obtained with the vacuum case, which is basically a Biot-Savart integral, shows that deviations of this order can be expected between measurements and simulations. Furthermore, this needs to be considered when determining the toroidal mode number, which depends in turn on a good charaterization of the poloidal one, as an error of $\delta m\pm 1$ leads to errors of $\delta n\pm 4$.
Beyond all possible hardware-related considerations or impact of surrounding structures on the measurements (non-ideal coils), it is essential to take all the previous issues into account whenever model validation against experimental results is attempted.
@@ Acknowledgements
The authors are grateful to K. Rahbarnia, R. Kleiber, S. Vaz Mendes, and C. Büschel for the helpful discussions and advice during the initial stages of this work. This work has been carried out within the framework of the EUROfusion Consortium, funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No 101052200 — EUROfusion). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them.
@@ Appendix A Physical model
The starting equations are Maxwell’s equations for a general strongly dispersive non-isotropic medium.
@@
×$\curl\mathbf{B}$ $=\mu_{0}\mathbf{J}_{\text{ext}}+\frac{1}{c^{2}}\frac{\partial\mathbf{E}}{\partial t}+\mu_{0}\mathbf{J}(\mathbf{E})$ (17)
@@ A.1 Vaccum electromagnetics
A.2 Plasma response
Plasma response is not included in the model described above since we have neglected the displacement currents. To evaluate the effect of the plasma on the radiated fields we start from the wave equation for $\mathbf{E}$, obtained by taking the curl of 16 and using 17,
@@ A.2 Plasma response
$\mathbf{k}\times\mathbf{k}\times\delta\mathbf{E}_{\mathbf{k},w}+\frac{\omega^{2}}{c^{2}}\epsilon_{\mathbf{k},\omega}\cdot\delta\mathbf{E}_{\mathbf{k},\omega}=-i\omega\mu_{0}\delta\mathbf{J}_{\mathbf{k},\omega}$ (33)
@@
$\delta\mathbf{B}(\mathbf{r},t)=\left[\int\delta\mathbf{B}_{\mathbf{k},w}\,d^{3}\mathbf{k}\right]e^{-i\omega t}$ (42)
@@
By evaluating the field at the positions of the coils we can calculate the synthetic signal and account for the response of an homogeneous plasma to an arbitrary distribution of currents. When investigating the effect of plasma response on the synthetic signals, and since the typical current distributions we are dealing with in this paper are far from being the ones that could be created in a homogeneous plasma, we will restrict ourselves to an ideal case that nevertheless maintains the essential characteristics of the problem. This is explained in detail in the main body of the paper.

### v3

@@ VI Conclusions and future work
One of the advantages of the tri-axial sensors of the helical arrays is that they allow us to study the polarization of the mode. The relationship between mode polarization and polarization of the measured radiation at the sensors position is not straightforward and the reconstruction of the mode polarization from the measurements needs prior knowledge of the measured mode numbers and a good measurement of the radial profile of the mode potentialPons-Villalonga et al. (2023) or, failing that, a synthetic database including several combinations of the former. This will be addressed in the near future using the experimental data presented inPons-Villalonga et al. (2023).
Finally, solving the wave equation by Fourier methods and including the response of an homogeneous plasma through its cold dielectric tensor, we have analyzed the plasma response for modes with $n=0$ and arbitrary mode number $m$. The model is limited in several important aspects, all of them discussed in appendix A.2, but it shows that measured mode numbers may differ from the real ones. Combining this result with the one obtained with the vacuum case, which is basically a Biot-Savart integral, shows that deviations of the order of $\Delta m=\pm 2$ or larger, even for modes not located in the plasma core, can be expected between measurements and simulations. Furthermore, this needs to be considered when determining the toroidal mode number, which depends in turn on a good charaterization of the poloidal one, as an error of $\delta m\pm 1$ leads to errors of $\delta n\pm 4$.
Beyond all possible hardware-related considerations or impact of surrounding structures on the measurements (non-ideal coils), it is essential to take all the previous issues into account whenever model validation against experimental results is attempted. We have taken an experimental case, in which part of the coils of the poloidal array were not working, to demonstrate the use of synthetic diagnostic both for mode number determination including noise modelling and polarization measurements.
The code for this synthetic diagnostic, that can be easily used to conduct similar analysis for other devices, is publicly available in the following GitHub repository: https://github.com/pponsv/synth_mirnov.
@@ Acknowledgements
The authors are grateful to K. Rahbarnia, R. Kleiber, S. Vaz Mendes, and C. Büschel for the helpful discussions and advice during the initial stages of this work. This work was partially funded by the Spanish Ministry of Science and Innovation under several contracts, Grant Nos. ENE2013-48109-P, FIS2017-88892-P, FIS2017-85252-R, PID2021-125607NB-I00, and by the European Regional Development Fund (ERDF) “A way of making Europe”. This work has been carried out within the framework of the EUROfusion Consortium, funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No 101052200 — EUROfusion). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them.
@@ Appendix A Physical model
The starting equations are Maxwell’s equations for a general strongly dispersive non-isotropic medium:
@@
×$\curl\mathbf{B}$ $=\mu_{0}\mathbf{J}_{\text{ext}}+\frac{1}{c^{2}}\frac{\partial\mathbf{E}}{\partial t}+\mu_{0}\mathbf{J}(\mathbf{E})\,,$ (17)
@@ A.1 Vaccum electromagnetics
A.2 Homogeneous plasma response for n=0 modes
All the results presented in the main body of the paper have been obtained using a Biot-Savart integral in vacuum, without taking into account the plasma response, which is not included in the model described above since we have neglected the displacement currents. To evaluate the effect of the plasma on the radiated fields we start from the wave equation for $\mathbf{E}$, obtained by taking the curl of 16 and using 17,
@@ A.2 Homogeneous plasma response for n=0 modes
$\mathbf{k}\times\mathbf{k}\times\delta\mathbf{E}_{\mathbf{k},\omega}+\frac{\omega^{2}}{c^{2}}\epsilon_{\mathbf{k},\omega}\cdot\delta\mathbf{E}_{\mathbf{k},\omega}=-i\omega\mu_{0}\delta\mathbf{J}_{\mathbf{k},\omega}$ (33)
@@
$\delta\mathbf{B}(\mathbf{r},t)=\left[\int\delta\mathbf{B}_{\mathbf{k},w}\,e^{i\mathbf{k}\cdot\mathbf{r}}d^{3}\mathbf{k}\right]e^{-i\omega t}$ (42)
Figure 28: On top, the full $n=0,m=3$ toroidally extended perturbation, with $s=0.5$ and $\sigma=0.03$, is shown. Bottom left shows the "straight" plasma column section. The same perturbation, radially localized around the detection plane of the poloidal array of Mirnov coils is shown on the bottom right.
@@
By evaluating the field at the positions of the coils we can calculate the synthetic signal and account for the effect that the response of a homogeneous plasma of a given density immersed in a constant magnetic field would have on an arbitrary distribution of currents. Since the typical current distributions we are dealing with in this paper are far from being the ones that could be created in a homogeneous plasma, we will restrict ourselves to an ideal case that nevertheless maintains the essential characteristics of the problem. In order to obtain a meaningful estimate of the plasma response while minimizing the impact of approximating the response of the real non-homogeneous plasma to that of a constant density one with magnetic field directed along the main direction of the real 3D field, we assume a highly localized mode current distribution centered toroidally at the toroidal plane where the poloidal array of Mirnov coils is located.
Both the extended and toroidally localized potential distributions are shown in figure 28. This approximation is valid as long as we restrict ourselves to $n=0$ modes. Only in this case the synthetic signals generated by the localized current distribution are a good approximation to the ones generated by a toroidally extended realistic distribution of currents. This is illustrated in figure 29, where the corresponding synthetic signals are compared. Only slight differences appear and the $m=3$ structure is well preserved in the case of the localized perturbation.
Figure 29: Synthetic signals generated by the toroidally extended (a) and localized (b) $n=0,m=3$ perturbation that is shown in figure 28.
To estimate the effect of the plasma response on the determination of the poloidal mode number, the localized distribution of currents, initially defined in the nodes of the boozer coordinates grid, is interpolated in a cartesian grid. For convenience, the grid and the current source are then rotated so that the $\mathbf{x}_{\text{cg}}$ direction of the cartesian grid is perpendicular to the plane of the poloidal array, which lies now in the plane formed by vectors $\mathbf{y}_{\text{cg}},\mathbf{z}_{\text{cg}}$ of the grid. The 3D Fourier transform of the current source in this grid, $\delta\mathbf{J}_{ext}(\mathbf{r})$, is computed numerically using a multidimensional FFT algorithm and used in equation 40. Since we want to highlight the effect of the plasma response, the field is only evaluated at those sensors of the poloidal array whose spacelike distribution most closely matches the shape of the plasma (see white dots in figures 31 and 32).
We first perform a sanity check without plasma response to ensure that the magnetic field calculated at the coil locations using the Biot-Savart integral over the cartesian grid and the solution 42 based on Fourier decomposition methods yield similar results both in phase and wave amplitude. Figure 30 shows the normalized synthetic signals obtained with both methods and the wave amplitudes measured by each coil. Also, figure 31 shows the real part of the components of the perturbed magnetic field calculated using the Fourier method, for $n=0,m=3$ and $n=0,m=9$ modes. The magnetic field perturbation that arises in response to the mode currents shows a variation of three orders of magnitude between the inner part of the plasma (milliTeslas) and the part where the coils are located (microTeslas) so we have chosen to represent the signed logarithm of the normalized values ($-\text{sign}(\delta b_{i})*\log(\absolutevalue{\delta b_{i}})$). This choice of representation emphasizes the structure of the magnetic field perturbation outside the plasma, which is our region of interest.
As discussed in the appendix, solving the equation in Fourier space allows us to describe the plasma response through the dispersion tensor of each plane wave component. By taking a constant density and a constant static field $B_{0}$ along $\mathbf{x}_{\text{cg}}$ in all the space domain we can take into account, approximately, the fast mobility of plasma electrons along the direction of the field and the subsequent shielding that they produce. Figure 32 shows the result of solving equation 31 considering now the contribution of the plasma response to the dispersion tensor given in 39. A homogeneous plasma with $n_{e}=10^{19}\ \text{m}^{-3}$, immersed in a $B_{0}=1$ T field, is taken to explore this effect.
Figure 31: Normalized components of the perturbed magnetic field in the central plane of the simulation grid $x_{\text{cg}}=0$ for $m=3$ (left) and $m=9$ (right). White circles indicate the location of the poloidal coils. By construction, the coils are sentitive only to $\delta B_{z}$ and $\delta B_{y}$.
Evolving in time the field perturbation, for each of the cases represented in figures 31 and 32, allows us to calculate the synthetic signals and apply the Lomb periodogram analysis. Figure 33 shows the the results of these analysis. As expected from the result anticipated in figure 30, the mode structure without plasma response is well diagnosed by the set of coils for all the mode numbers considered. On the contrary, the plasma response clearly modifies the field structure and consequently the mode number detected by the coils differs from the input ones, the differences $\Delta m\equiv m_{out}-m_{in}$ reaching values up to $\Delta m=\pm 2$. This can be visually observed in figures 31 and 32 where it is shown that the structure of maxima and minima of $\delta\mathbf{B}$ is modified by the response of the plasma, in particular for those components of the magnetic field that have an impact on the coils measurement ($\delta B_{z}$ and $\delta B_{y}$). Note that the component of the perturbed field along the direction of the static homogeneous field ($\delta B_{x}$), which is not detected by the coils, is precisely the component less modified by the plasma response since mode currents flow mainly in the $\mathbf{x}_{\text{cg}}$ direction and thus barely contribute to changes in $\delta B_{x}$.
The approximations made in the plasma response model (i.e. modes with n=0 and toroidally localized current source), do not allow us to estimate with certainty the changes that the plasma response induces in the measured mode number. Indeed, the detected mode numbers show differences with respect to the initial ones. However, this can only serve as an indication that the plasma has a clear effect on mode detection, but we cannot provide a clear quantification of this since the price of assuming an homogeneous plasma is that the electric field associated with the mode produces currents in the whole volume and not only in the part where the plasma is actually present. This can only be addressed with a full-wave code that solves this problem in real space.
@@ Appendix B Phase delay term for the 3D Lomb periodogram
Appendix C Wave field attenuation
We inquire now about the effects of coil distance over the measured signal. For that, a set of signals with different mode numbers have been simulated for a hypothetical linear array of coils located at $\varphi=0$, $z=0$ and distributed with increasing radial distance $R$. The amplitude of the perturbed field at the $i$-th coil position is
$A_{i}\equiv\sqrt{\Bigl{\langle}\absolutevalue{\delta\dot{\mathbf{B}_{i}}}^{2}\Bigr{\rangle}_{t}}\,,$ (49)
where $\langle\cdot\rangle_{t}$ denotes a time average. In the multipolar approximation Jackson (1999), this signal follows an inverse-power law, decaying as $1/r^{m+1}$, so an attempt has been made to fit the field amplitudes to the function
$A(r)=\dfrac{A_{0}(m)}{(r_{\text{LCFS}}+\delta r)^{m+1}}\,,$ (50)
where $A_{0}(m)$ is a mode-dependent initial amplitude, $r_{\text{LCFS}}$ is the distance between the coil and the LCFS, and $\delta r$ is a free parameter. Figure 34 shows the results of the fit of this function to the simulated perturbation amplitudes for a set of $n=0$ modes. For the fit, while $A_{0}$ is $m$-dependant and thus is fitted separately for each value of $m$, $\delta r$ is fitted simultaneously to all signals, ensuring it remains constant. As the multi-polar approximation is valid only when the measuring point is reasonably far from the plasma, the first coils have been omitted from the fit. The goodness of the fit for these few first coils improves with increasing mode number, but for $m=1$ notable deviations can be observed. The obtained value of $\delta r$ is consistent with a multipole located approximately over the central field conductor of the device, instead of being over the magnetic axis as expected. This is probably due to the shape of the plasma column, that differs notably from a cylinder, so some deviation from the ideal case is to be expected. As a diagnostic, an arrangement like this would be of little use, as the contributions from the toroidal and poloidal modes would be mixed and very difficult to decouple, and would require a space between the LCFS and the first wall that in TJ-II is not available.
Appendix D Performance
As the time evolution of the mode can be decoupled from the Biot-Savart integral over the plasma, the code is very performant. For a $n_{s}\times n_{\theta}\times n_{\varphi}=100\times 150\times 1000$ grid, the integration of a single mode takes around 5 seconds per coil on an 8-core desktop computer. The code is mainly written in Fortran, making use of the FFTW3 library Frigo and Johnson (2005) for the calculation of fast Fourier transforms and is parallelized using OpenMP; it also has a Python API for ease of use.